# The Ranking-Selberg integral on ${\bf GSp}(2)$ for square free levels

### v1

@@ 6.1. Local zeta integral for Bessel models
(6.23) $\begin{array}[]{|c|c|c|c|c|c|c|c|}\hline\cr&{{\rm type}}&{\phi}&{B}&{\mu}&L/F&{g}\\ \hline\cr{\rm case\,1}&{\text{I or IIb}}&{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}}&{K{\rm-fixed}}&{{\rm unramified}}&-&{1_{4}}\\ \hline\cr{\rm case\,2}&{\text{I or IIb}}&{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}^{e}\mathfrak{o}_{L}\oplus 1+{\mathfrak{p}}^{e}\mathfrak{o}_{L}}}&{K{\rm-fixed}}&{{\rm ramified}}&{\rm inert}&{b}\\ \hline\cr{\rm case\,3}&{\text{I or IIb}}&{\widehat{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}^{e}\mathfrak{o}_{L}\oplus 1+{\mathfrak{p}}^{e}\mathfrak{o}_{L}}}}&{K{\rm-fixed}}&{{\rm ramified}}&{\rm inert}&{b}\\ \hline\cr{\rm case\,4}&{\text{I or IIb}}&{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}}&{K_{0}({\mathfrak{p}}){\rm-fixed}}&{{\rm unramified}}&{\rm inert}&{\eta}\\ \hline\cr{\rm case\,5}&{\text{IIIa}}&{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}}&{K_{0}({\mathfrak{p}}){\rm-fixed}}&{{\rm unramified}}&{\rm inert}&{\eta}\\ \hline\cr{\rm case\,6}&{\text{VIb}}&{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}}&{K_{0}({\mathfrak{p}}){\rm-fixed}}&{{\rm unramified}}&{\rm inert}&{\eta}\\ \hline\cr\end{array}$
Here,
(6.24) $b:=\left[\begin{smallmatrix}\varpi^{e}1_{2}&S^{\dagger}\\ 0&1_{2}\end{smallmatrix}\right]\in\mathsf{G}(F),$
@@
As in [32], for smooth representation $(\pi,V_{\pi})$ of $\mathsf{G}(F)$, we define an operator $T_{m,n}\in{\operatorname{End}}(V_{\pi})$ by
$T_{m,n}v:={\operatorname{vol}}(K_{0}({\mathfrak{p}}))\int_{K_{0}({\mathfrak{p}})h(l,m)K_{0}({\mathfrak{p}})}\pi(g)v\,{\rm{d}}g,\quad v\in V_{\pi}$
@@
If $\pi$ is irreducible and has the $(S,\Lambda)$-Bessel models, then for any $B\in{\mathscr{B}}(S,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$, by [32, Lemma 4.4 (i)], we have the support conditions:
@@ 6.2. Unramified computation (case1)
When $\pi$ is spherical, $\pi$ is of type I or IIb in our cases. Recall that there exists a unique element $B_{\pi}^{0}\in{\mathscr{B}}(S,\Lambda)[\pi]^{K}$ such that $B_{\pi}^{0}(1_{4})=1$. Then, the following formula for an unramified $\mu$ in its greatest generality is due to Sugano ([43, Tehorem 2.1]):
@@
In [20, Proposition 5.9], a proof of this formula is given when $L=F\oplus L$ and $\pi$ is of type I based on the explicit formula of $B_{\pi}^{0}$ due to [6].
6.3. Unramified computation (case2 & case3)
Now we assume that $e:={\rm cond}(\mu)>0$ and that $S\in{\rm{Sym^{2}}}(F)$ satisfies the following conditions;
@@ 6.3. Unramified computation (case2 & case3)
  • $S\in{\bf GL}_{2}(\mathfrak{o}),\quad-\frac{{\mathsf{d}}}{2}={\operatorname{tr}}(SS^{\dagger})\in\mathfrak{o}^{\times}.$
Then $\varpi$ remains a prime element in ${\mathfrak{p}}\mathfrak{o}_{L}$. Let $b$ as in (6.24), and set $\phi=\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}^{e}\mathfrak{o}_{L}\oplus 1+{\mathfrak{p}}^{e}\mathfrak{o}_{L}}\in\mathcal{S}(L^{2})$. Our goal of this subsection is to calculate the zeta integrals $Z(\phi,B_{\pi}^{0},s,\mu;b)$ and $Z(\widehat{\phi},B_{\pi}^{0},s,\mu;b)$ explicitly. We remark that $Z(\phi,B_{\pi}^{0},s,\mu;1_{4})$ must be zero because the $a$-integral in (6.8) vanishes.
@@
Note that this sum has the absolute value $1$ and is independent of the choice of $\varpi$. The following lemma is more or less well-known ([45, (2.18)]).
@@
For a virtual representation $V$ of the Weil group $\mathfrak{W}_{F}$ of $F$, let $\varepsilon(s,V,\psi)$ denote the local epsilon factor a la Deligne-Langlands ([46]). By ([46, (3.4.8)]), we have
@@
$f_{\phi}^{(s,\Lambda,\mu)}(k^{\#})$ =$=\begin{cases}q^{-2e+2}(q^{2}-1)^{-1}\mu_{L}(d)^{-1}\mu(\det k^{\#})\quad&(c\in{\mathfrak{p}}^{e}\mathfrak{o}_{L},d\in 1+{\mathfrak{p}}^{e}\mathfrak{o}_{L}),\\ 0&({\rm otherwise}),\end{cases}$
@@
First, we prove the first equation. By a direct calculation, we have
$f_{\phi}^{(s,\Lambda,\mu)}(k^{\#})$ $=\mu(\det k^{\#})\int_{L^{\times}}\phi([\begin{smallmatrix}c\tau^{\dagger}&d\tau^{\dagger}\end{smallmatrix}])\Lambda\mu_{L}(\tau)|\tau\tau^{\dagger}|^{s+1}{\rm{d}}^{\times}\tau$
$=\mu(\det k^{\#})\int_{L^{\times}}\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}^{e}\mathfrak{o}_{L}}(c\tau^{\dagger})\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{1+{\mathfrak{p}}^{e}\mathfrak{o}_{L}}(d\tau^{\dagger})\Lambda\mu_{L}(\tau)|\tau\tau^{\dagger}|^{s+1}{\rm{d}}^{\times}\tau.$
Since $c\in\mathfrak{o}_{L}^{\times}$ or $d\in\mathfrak{o}_{L}^{\times}$, it is easy to check that $(c\tau^{\dagger},d\tau^{\dagger})\in{\mathfrak{p}}^{e}\mathfrak{o}_{L}\times(1+{\mathfrak{p}}^{e}\mathfrak{o}_{L})$ if and only if
$c\in{\mathfrak{p}}^{e}\mathfrak{o}_{L},\quad d\in\mathfrak{o}_{L}^{\times},\quad\tau\in(d^{\dagger})^{-1}(1+{\mathfrak{p}}^{e}\mathfrak{o}_{L}).$
Hence, if $c\in{\mathfrak{p}}^{e}\mathfrak{o}_{L}$, $d\in\mathfrak{o}_{L}^{\times}$, we get
$f_{\phi}^{(s,\Lambda,\mu)}(k^{\#})$ $=\mu(\det k^{\#})\cdot{\operatorname{vol}}(1+{\mathfrak{p}}^{e}\mathfrak{o}_{L})\cdot\Lambda\mu_{L}((d^{\dagger})^{-1})$
$=q^{-2e+2}(q^{2}-1)^{-1}\mu_{L}(d)^{-1}\mu(\det k^{\#}).$
Next, we prove the second equation. It is easy to check that
$\widehat{\phi}(x,y)=q^{-4e}\psi_{L}(x)\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}^{{}^{e}}\mathfrak{o}}(x)\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}^{{}^{e}}\mathfrak{o}}(y),\quad x,y\in L.$
Hence we have
$f_{\widehat{\phi}}^{(s,\Lambda,\mu)}(k^{\#})$ $=\mu(\det k^{\#})\int_{L^{\times}}\widehat{\phi}([\begin{smallmatrix}c\tau^{\dagger}&d\tau^{\dagger}\end{smallmatrix}])\Lambda\mu_{L}(\tau)|\tau\tau^{\dagger}|_{p}^{s+1}{\rm{d}}^{\times}\tau$
$=q^{-4e}\mu(\det k^{\#})\int_{L^{\times}}\psi_{L}(c\tau^{\dagger})\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}^{-e}\mathfrak{o}_{L}}(c\tau^{\dagger})\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}^{-e}\mathfrak{o}_{L}}(d\tau^{\dagger})\Lambda\mu_{L}(\tau)|\tau\tau^{\dagger}|_{p}^{s+1}{\rm{d}}^{\times}\tau$
(6.28) $=q^{-4e}\mu(\det k^{\#})\sum_{n{\geqslant}-e}\Lambda\mu_{L}(\varpi)^{n}q^{-2n(s+1)}\int_{\mathfrak{o}_{L}^{\times}}\psi_{L}(c\varpi^{n}\tau)\mu_{L}(\tau){\rm{d}}^{\times}\tau.$
We put $c=c_{0}\varpi^{m}$, $c_{0}\in\mathfrak{o}_{L}^{\times}$. Then by Lemma 6.1, with $(F,\mu,\psi)$ replaced by $(L,\Lambda\mu_{L},\psi_{L})$, we have
$\int_{\mathfrak{o}_{L}^{\times}}\psi_{L}(c\varpi^{n}\tau)\mu_{L}(\tau){\rm{d}}^{\times}\tau$ $=\mu_{L}(c_{0})^{-1}\int_{\mathfrak{o}_{L}^{\times}}\psi_{L}(\varpi^{n+m}\tau)\mu_{L}(\tau){\rm{d}}^{\times}\tau$
$=\delta_{n,-e-m}q^{-e+2}(q^{2}-1)^{-1}\mu_{L}(c_{0}^{-1}\varpi)^{e}W_{L}(\mu_{L},\psi_{L}).$
[…]
$f_{\widehat{\phi}}^{(s,\Lambda,\mu)}(k^{\#})$ $=q^{-5e+2}(q^{2}-1)^{-1}\Lambda(\varpi)^{-e}\mu_{L}(c_{0})^{-1}W_{L}(\mu_{L},\psi_{L})\sum_{n{\geqslant}-e}\delta_{n,-e-m}\Lambda\mu_{L}(\varpi)^{n+e}q^{-2n(s+1)}$
=$=\begin{cases}q^{e(2s-3)+2}(q^{2}-1)^{-1}\Lambda(\varpi)^{-e}\mu_{L}(c)^{-1}W_{L}(\mu_{L},\psi_{L})\quad&(m=0),\\ 0&(m>0).\end{cases}\qed$
@@
$\sum_{\begin{subarray}{c}\eta\in\mathfrak{o}_{L}/{\mathfrak{p}}^{e}\mathfrak{o}_{L}\\ u+{\operatorname{N}}_{L/F}(\eta)\in\mathfrak{o}^{\times}\end{subarray}}\mu(u+{\operatorname{N}}_{L/F}(\eta))=(-1)^{e}q^{e}\mu(u).$
@@
Because $L/F$ is an unramified extension, the norm map $\mathfrak{o}_{L}^{\times}\owns\eta\mapsto{\operatorname{N}}_{L/F}(\eta)\in\mathfrak{o}^{\times}$ is surjective. Thus there exists $v\in\mathfrak{o}_{L}^{\times}$ such that $u={\operatorname{N}}_{L/F}(v)$. By replacing $\eta$ by $v\eta$, without loss of generality, we can assume that $u=1$.
@@
$\overline{\mathfrak{o}_{L,i}}=\begin{cases}\left\{\eta\in\left(\mathfrak{o}_{L}/{\mathfrak{p}}^{e}\mathfrak{o}_{L}\right)^{\times}\ \middle|\ 1+{\operatorname{N}}_{L/F}(\eta)\not\equiv 0\ (\text{mod }{\mathfrak{p}})\right\}\quad&(i=0),\\ {\mathfrak{p}}^{i}\mathfrak{o}_{L}/{\mathfrak{p}}^{e}\mathfrak{o}_{L}&(0<i{\leqslant}e),\\ \varnothing&(i=e+1)\end{cases}$
@@
$\overline{\mathfrak{o}_{L,i}}\setminus\overline{\mathfrak{o}_{L,i+1}}\owns\eta\ (\text{mod }{\mathfrak{p}}^{e}\mathfrak{o}_{L})\mapsto 1+N_{L/F}(\eta)\ (\text{mod }{\mathfrak{p}}^{e})\in\overline{{\mathcal{U}}_{2i}}\setminus\overline{{\mathcal{U}}_{2i+1}}$
@@
$\sum_{\begin{subarray}{c}\eta\in\mathfrak{o}_{L}/{\mathfrak{p}}^{e}\mathfrak{o}_{L}\\ 1+{\operatorname{N}}_{L/F}(\eta)\in\mathfrak{o}^{\times}\end{subarray}}\mu(1+{\operatorname{N}}_{L/F}(\eta))$ $=\sum_{i=0}^{e}\sum_{\eta\in\overline{\mathfrak{o}_{L,i}}\setminus\overline{\mathfrak{o}_{L,i+1}}}\mu(1+{\operatorname{N}}_{L/F}(\eta))$
@@
Let $B\in{\mathscr{B}}(S,\Lambda)[\pi]^{K}$ and $\eta\in\mathfrak{o}$ with $\frac{{\mathsf{a}}^{6}{\mathsf{d}}}{4}+{\operatorname{N}}_{L/F}(\eta)\in\varpi^{j}\mathfrak{o}^{\times}$ $(0{\leqslant}j{\leqslant}e)$.
@@
    If we set $Y_{\eta}=-{\mathsf{a}}^{2}S^{\dagger}+X_{\eta}$, then we have
@@
(i) By noting that $\det S=-{\mathsf{d}}/4$, $\det X_{\eta}=-{\mathsf{a}}^{-2}{\operatorname{N}}_{L/F}(\eta)$, and ${\operatorname{tr}}(SX_{\eta})={\operatorname{tr}}(X_{\eta}S^{\dagger})=0$, we obtain
$\det Y_{\eta}=\tfrac{1}{2}{\operatorname{tr}}(Y_{\eta}Y_{\eta}^{\dagger})=\tfrac{1}{2}{\operatorname{tr}}({\mathsf{a}}^{4}SS^{\dagger}+X_{\eta}^{\dagger}X_{\eta})={\mathsf{a}}^{4}\det S+\det(X_{\eta})=-\tfrac{{\mathsf{a}}^{4}{\mathsf{d}}}{4}-{\mathsf{a}}^{-2}{\operatorname{N}}_{L/F}(\eta).$
(ii) By (i) and the assumption that $\frac{{\mathsf{a}}^{6}{\mathsf{d}}}{4}+{\operatorname{N}}_{L/F}(\eta)\in\varpi^{j}\mathfrak{o}^{\times}$, we have $Y_{\eta}\in{\bf GL}_{2}(F)$. The disjoint union
${\bf GL}_{2}(F)=\sqcup_{m{\geqslant}0}L^{\times}\left[\begin{smallmatrix}\varpi^{m}&0\\ 0&1\end{smallmatrix}\right]{\bf GL}_{2}(\mathfrak{o}_{L}).$
from [43, Lemma 2–4] implies that there exist $m\in{\mathbb{Z}}_{{\geqslant}0}$, $\ell\in{\mathbb{Z}}$, $\tau\in\mathfrak{o}_{L}^{\times}$, $A\in{\bf GL}_{2}(\mathfrak{o})$, such that
@@
We now prove that $m=j$ and $\ell=0$. Because $\tau\in{\bf GL}_{2}(\mathfrak{o})$, the Smith normal form of $Y_{\eta}$ is $\left[\begin{smallmatrix}\varpi^{m+\ell}&0\\ 0&\varpi^{\ell}\end{smallmatrix}\right]$. Hence by the theory of the Smith normal form, we only have to show that the elements of $Y_{\eta}$ are coprime. This follows from the identity ${\operatorname{tr}}(Y_{\eta}S)=\frac{{\mathsf{a}}^{2}{\mathsf{d}}}{2}$ and the assumption $\frac{{\mathsf{d}}}{2}\in\mathfrak{o}^{\times}$.
@@
$\iota(\left[\begin{smallmatrix}1&0\\ -\eta^{\dagger}&1\end{smallmatrix}\right])=\left[\begin{smallmatrix}1_{2}&0\\ X_{\eta}^{\dagger}&1_{2}\end{smallmatrix}\right],\quad\iota(\left[\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\right])b\iota(\left[\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\right])=\left[\begin{smallmatrix}1_{2}&0\\ -{\mathsf{a}}^{2}S&\varpi^{e}1_{2}\end{smallmatrix}\right].$

### v2

@@ 6.1. Local zeta integral for Bessel models
(6.19) $\begin{array}[]{|c|c|c|c|c|c|c|c|c|}\hline\cr&{{\rm type}}&{\phi}&{B}&{\Lambda}&{\mu}&L/F&{g}\\ \hline\cr{\rm case\,1}&{\text{I or IIb}}&{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}}&{K{\rm-fixed}}&{{\rm unramified}}&{{\rm unramified}}&-&{1_{4}}\\ \hline\cr{\rm case\,2}&{\text{I or IIb}}&{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}_{L}^{e}\oplus 1+{\mathfrak{p}}_{L}^{e}}}&{K{\rm-fixed}}&{{\rm unramified}}&{{\rm ramified}}&{\rm inert}&{b}\\ \hline\cr{\rm case\,3}&{\text{I or IIb}}&{\widehat{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}_{L}^{e}\oplus 1+{\mathfrak{p}}_{L}^{e}}}}&{K{\rm-fixed}}&{{\rm unramified}}&{{\rm ramified}}&{\rm inert}&{b}\\ \hline\cr{\rm case\,4}&{\text{I or IIb}}&{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}_{L}^{f}\oplus 1+{\mathfrak{p}}_{L}^{f}}}&{K{\rm-fixed}}&{{\rm ramified}}&{{\rm unramified}}&{\rm inert}&{b^{\prime}}\\ \hline\cr{\rm case\,5}&{\text{I or IIb}}&{\widehat{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}_{L}^{f}\oplus 1+{\mathfrak{p}}_{L}^{f}}}}&{K{\rm-fixed}}&{{\rm ramified}}&{{\rm unramified}}&{\rm inert}&{b^{\prime}}\\ \hline\cr{\rm case\,6}&{\text{I or IIb}}&{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}}&{K_{0}({\mathfrak{p}}){\rm-fixed}}&{{\rm unramified}}&{{\rm unramified}}&{\rm inert}&{\eta}\\ \hline\cr{\rm case\,7}&{\text{IIIa}}&{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}}&{K_{0}({\mathfrak{p}}){\rm-fixed}}&{{\rm unramified}}&{{\rm unramified}}&{\rm inert}&{\eta}\\ \hline\cr{\rm case\,8}&{\text{VIb}}&{\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{L}\oplus\mathfrak{o}_{L}}}&{K_{0}({\mathfrak{p}}){\rm-fixed}}&{{\rm unramified}}&{{\rm unramified}}&{\rm inert}&{\eta}\\ \hline\cr\end{array}$
Here, the integer $e$ (resp. $f$) is the conductor of $\mu$ (resp. $\Lambda$) defined to be the minimum non-negative integer $n$ satisfying $\mu|_{(1+{\mathfrak{p}}^{n})\cap{\mathfrak{o}}^{\times}}=1$ (resp. $\Lambda|_{(1+{\mathfrak{p}}_{L}^{n})\cap{\mathfrak{o}}_{L}^{\times}}=1$), $b,b^{\prime}\in\mathsf{G}(F)$ are given as
(6.20) $b:=\left[\begin{smallmatrix}\varpi^{e}1_{2}&T^{\dagger}\\ 0&1_{2}\end{smallmatrix}\right],\quad b^{\prime}:=\left[\begin{smallmatrix}\varpi^{2f}&&&\\ &\varpi^{f}&&\\ &&1&\\ &&&\varpi^{f}\end{smallmatrix}\right],$
@@
As in [33], for smooth representation $(\pi,V_{\pi})$ of $\mathsf{G}(F)$, we define an operator $T_{\ell,m}\in{\operatorname{End}}(V_{\pi})$ by
$T_{\ell,m}v:={\operatorname{vol}}(K_{0}({\mathfrak{p}}))\int_{K_{0}({\mathfrak{p}})h(\ell,m)K_{0}({\mathfrak{p}})}\pi(g)v\,{\rm{d}}g,\quad v\in V_{\pi}$
@@
If $\pi$ is irreducible and has the $(T,\Lambda)$-Bessel models, then for any $B\in{\mathscr{B}}(T,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$, by [33, Lemma 4.4 (i)], we have the support conditions:
@@ 6.2. Unramified computation (case 1)
When $\pi$ is spherical, $\pi$ is of type I or IIb in our cases. Recall that there exists a unique element $B_{\pi}^{0}\in{\mathscr{B}}(T,\Lambda)[\pi]^{K}$ such that $B_{\pi}^{0}(1_{4})=1$. Then, the following formula for an unramified $\mu$ in its greatest generality is due to Sugano ([44, Theorem 2.1]):
@@
In [21, Proposition 5.9], a proof of this formula is given when $L=F\oplus F$ and $\pi$ is of type I based on the explicit formula of $B_{\pi}^{0}$ due to [6].
6.3. Unramified computation (case 2 and case 3)
Now we assume that $e>0$ and that $T\in{\rm{Sym^{2}}}(F)$ satisfies the following conditions;
@@ 6.3. Unramified computation (case 2 and case 3)
  • $T\in{\bf GL}_{2}(\mathfrak{o}),\quad-\frac{{\mathsf{d}}}{2}={\operatorname{tr}}(TT^{\dagger})\in\mathfrak{o}^{\times}.$
Then $\varpi$ remains a prime element in ${\mathfrak{p}}_{L}$. Let $b$ be as in (6.20), and set $\phi=\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}_{L}^{e}\oplus 1+{\mathfrak{p}}_{L}^{e}}\in\mathcal{S}(L^{2})$. Our goal of this subsection is to calculate the zeta integrals $Z(\phi,B_{\pi}^{0},s,\mu;b)$ and $Z(\widehat{\phi},B_{\pi}^{0},s,\mu;b)$ explicitly. We remark that $Z(\phi,B_{\pi}^{0},s,\mu;1_{4})$ must be zero because the $a$-integral in (6.8) vanishes.
@@
Note that this sum has the absolute value $1$ and is independent of the choice of $\varpi$. The following lemma is more or less well-known ([46, (2.18)]).
@@
For a virtual representation $V$ of the Weil group $\mathfrak{W}_{F}$ of $F$, let $\varepsilon(s,V,\psi)$ denote the local epsilon factor a la Deligne-Langlands ([47]). By ([47, (3.4.8)]), we have
@@
$f_{\phi}^{(s,\Lambda,\mu)}(k^{\#})$ =$=\begin{cases}q^{-2e+2}(q^{2}-1)^{-1}\mu_{L}(d)^{-1}\mu(\det k^{\#})\quad&(c\in{\mathfrak{p}}_{L}^{e}),\\ 0&({\rm otherwise}),\end{cases}$
@@
These equations immediately follow by Lemma 6.1. ∎
@@
$\sum_{\begin{subarray}{c}\eta\in\mathfrak{o}_{L}/{\mathfrak{p}}_{L}^{e}\\ u+{\operatorname{N}}_{L/F}(\eta)\in\mathfrak{o}^{\times}\end{subarray}}\mu(u+{\operatorname{N}}_{L/F}(\eta))=(-1)^{e}q^{e}\mu(u).$
@@
Because $L/F$ is an unramified extension, the norm map ${\operatorname{N}}_{L/F}:\mathfrak{o}_{L}^{\times}\rightarrow\mathfrak{o}^{\times}$ is surjective. Thus, there exists $v\in\mathfrak{o}_{L}^{\times}$ such that $u={\operatorname{N}}_{L/F}(v)$. By replacing $\eta$ by $v\eta$, without loss of generality, we may assume that $u=1$.
@@
$\overline{\mathfrak{o}_{L,i}}=\begin{cases}\left\{\eta\in\left(\mathfrak{o}_{L}/{\mathfrak{p}}_{L}^{e}\right)^{\times}\ \middle|\ 1+{\operatorname{N}}_{L/F}(\eta)\not\equiv 0\ (\text{mod }{\mathfrak{p}})\right\}\quad&(i=0),\\ {\mathfrak{p}}_{L}^{i}/{\mathfrak{p}}_{L}^{e}&(0<i{\leqslant}e),\\ \varnothing&(i=e+1)\end{cases}$
@@
$\overline{\mathfrak{o}_{L,i}}\setminus\overline{\mathfrak{o}_{L,i+1}}\owns\eta\ (\text{mod }{\mathfrak{p}}_{L}^{e})\mapsto 1+N_{L/F}(\eta)\ (\text{mod }{\mathfrak{p}}^{e})\in\overline{{\mathcal{U}}_{2i}}\setminus\overline{{\mathcal{U}}_{2i+1}}$
@@
$\sum_{\begin{subarray}{c}\eta\in\mathfrak{o}_{L}/{\mathfrak{p}}_{L}^{e}\\ 1+{\operatorname{N}}_{L/F}(\eta)\in\mathfrak{o}^{\times}\end{subarray}}\mu(1+{\operatorname{N}}_{L/F}(\eta))$ $=\sum_{i=0}^{e}\sum_{\eta\in\overline{\mathfrak{o}_{L,i}}\setminus\overline{\mathfrak{o}_{L,i+1}}}\mu(1+{\operatorname{N}}_{L/F}(\eta))$
@@
Let $B\in{\mathscr{B}}(T,\Lambda)[\pi]^{K}$ and $\eta\in\mathfrak{o}$ with $\frac{{\mathsf{a}}^{6}{\mathsf{d}}}{4}+{\operatorname{N}}_{L/F}(\eta)\in\varpi^{j}\mathfrak{o}^{\times}$ $(0{\leqslant}j{\leqslant}e)$.
@@
    If we set $Y_{\eta}=-{\mathsf{a}}^{2}T^{\dagger}+X_{\eta}$, then we have
@@
(i) By noting that $\det T=-{\mathsf{d}}/4$, $\det X_{\eta}=-{\mathsf{a}}^{-2}{\operatorname{N}}_{L/F}(\eta)$, and ${\operatorname{tr}}(TX_{\eta})={\operatorname{tr}}(X_{\eta}T^{\dagger})=0$, we obtain
$\det Y_{\eta}=\tfrac{1}{2}{\operatorname{tr}}(Y_{\eta}Y_{\eta}^{\dagger})=\tfrac{1}{2}{\operatorname{tr}}({\mathsf{a}}^{4}TT^{\dagger}+X_{\eta}^{\dagger}X_{\eta})={\mathsf{a}}^{4}\det T+\det(X_{\eta})=-\tfrac{{\mathsf{a}}^{4}{\mathsf{d}}}{4}-{\mathsf{a}}^{-2}{\operatorname{N}}_{L/F}(\eta).$
(ii) By (i) and the assumption that $\frac{{\mathsf{a}}^{6}{\mathsf{d}}}{4}+{\operatorname{N}}_{L/F}(\eta)\in\varpi^{j}\mathfrak{o}^{\times}$, we have $Y_{\eta}\in{\bf GL}_{2}(F)$. The disjoint union ${\bf GL}_{2}(F)=\sqcup_{m{\geqslant}0}L^{\times}\left[\begin{smallmatrix}\varpi^{m}&0\\ 0&1\end{smallmatrix}\right]{\bf GL}_{2}(\mathfrak{o})$ ([44, Lemma 2–4]) implies that there exist $m\in{\mathbb{Z}}_{{\geqslant}0}$, $\ell\in{\mathbb{Z}}$, $\tau\in\mathfrak{o}_{L}^{\times}$, $A\in{\bf GL}_{2}(\mathfrak{o})$, such that
@@
We now prove that $m=j$ and $\ell=0$. Because $\tau\in{\bf GL}_{2}(\mathfrak{o})$, the Smith normal form of $Y_{\eta}$ is $\left[\begin{smallmatrix}\varpi^{m+\ell}&0\\ 0&\varpi^{\ell}\end{smallmatrix}\right]$. Hence by the theory of the Smith normal form, we only have to show that the elements of $Y_{\eta}$ are coprime. This follows from the identity ${\operatorname{tr}}(Y_{\eta}T)=\frac{{\mathsf{a}}^{2}{\mathsf{d}}}{2}$ and the assumption $\frac{{\mathsf{d}}}{2}\in\mathfrak{o}^{\times}$.
@@
$\iota(\left[\begin{smallmatrix}1&0\\ -\eta^{\dagger}&1\end{smallmatrix}\right])=\left[\begin{smallmatrix}1_{2}&0\\ X_{\eta}^{\dagger}&1_{2}\end{smallmatrix}\right],\quad\iota(\left[\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\right])b\iota(\left[\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\right])=\left[\begin{smallmatrix}1_{2}&0\\ -{\mathsf{a}}^{2}T&\varpi^{e}1_{2}\end{smallmatrix}\right].$